# IsoSEL: Isometric Structural Entropy Learning for Deep Graph Clustering in Hyperbolic Space

### v1

@@ VIII-A1 Datasets & Baselines
VGAE ARGA MVGRL Sublime GRACE DCRN gCooL S3GC DinkNet GC-Flow DGCluster IsoSEL Avg. Gain
Cora 64.72 68.40 80.03 71.30 73.87 66.91 77.03 74.27 78.11 78.92 76.45 81.77 7.22
Citeseer 61.00 60.72 61.23 68.53 63.12 70.86 71.35 68.80 70.36 70.06 70.14 71.37 6.26
AMAP 74.26 69.28 41.07 47.22 67.66 77.94 68.52 75.15 79.23 72.54 78.92 78.68 7.51
Computer 42.44 45.67 29.53 45.52 46.62 30.42 44.50 39.03 43.52 43.20 43.65 47.34 5.14
Pubmed 68.31 64.82 67.01 66.62 63.72 69.63 68.24 71.30 69.42 73.42 66.39 73.53 6.26
TABLE III: Comparison of time complexity and time cost for baselines.
Method Time Complexity Inference Time (s)
DGI $\mathcal{O}(ED+Nd^{2})$ 20.15
MVGRL $\mathcal{O}(N^{2}d+Nd^{2})$ 176.35
S3GC $\mathcal{O}(NSd^{2})$ 520.41
Dink-Net $\mathcal{O}(BKd+K^{2}d+Bd)$ 37.32
DGCluster $\mathcal{O}(K^{2}N+P^{2}N)$ 8.56
IsoSEL (Ours) $\mathcal{O}(N+mN)$ 7.53
@@ VIII-A3 Model Configuration & Reproducibility
We instantiate the proposed IsoSEL in the Lorentz model of hyperbolic space with a standard curvature of $\kappa=-1$. The model configuration by default is specified as follows. We consider the partitioning tree of height $2$ by default and hyperparameter $N_{1}$ of network architecture is set as $10$. The augmentation ratio $\alpha$ is $0.2$. We stack the Lorentz convolution once in LSEnet, and set the number of negative samples to $30$ in Lorentz tree contrastive learning. The parameters are optimized by Adam and further details are provided in implementation note of Appendix VI. As for the baselines, the hyperparameter setting is the same as that of original papers.
@@ VIII-B1 Effectiveness
We report the clustering results in terms of NMI and ARI on the five real datasets in Table I, and ACC in Table II. Deep clustering models require cluster number $K$ in training phase, except DGCluster, RGC and our IsoSEL. For clustering models without $K$, DGCluster conducts self-supervised representation learning with modularity [44], and RGC searches the optimal $K$ with reinforcement learning [8]. For Riemannian models, i.e., Congregate and IsoSEL, we project Euclidean attributes onto the respective manifold, and the projection is given by exponential map with the reference point of the origin. DCRN does not scale well to the larger Computer dataset, and runs into Out-Of-Memory (OOM) error on our hardware. In Table I, the proposed IsoSEL has the best clustering results except two cases, NMI on AMAP and ARI on Computer. Note that, IsoSEL outperforms $14$ recent deep clustering models by $+1.3\%$ NMI on average, and generally achieves the best ACC in Table II. Also, we observe that IsoSEL shows more consistent performance over different datasets, e.g., DGCluster obtain the best NMI on AMAP but is unsatisfactory on Computer.
@@ VIII-B2 Efficiency
Table III collects the theoretical time complexity and empirical inference time on Cora. The inference time is given in seconds, and we are presented as the fastest deep model in the table. As for theoretical complexity, $N$ and $E$ denote the number of nodes and edges, respectively. $B$ is the batch size of DinkNet. $S$ is the average degree of nodes and thereby the complexity of S3GC is related to the number of edges. In DGCluster, $P$ is a dimensionality of auxiliary matrix, $P\ll N$. In IsoSEL, $m$ is the number of negative samples, and it yields the linear complexity thanks to the reformulation of differential structural information and hyperbolic reweighing of Lorentz tree contrastive learning. It is noteworthy to mention that, compared to recent deep clustering model such as S3GC, we achieve superior clustering effectiveness while removing the prior knowledge of cluster number without computational overhead. The following parts will demonstrate how the differential formulation, hyperbolic space and Lorentz tree contrastive learning contribute to the success of the proposed model.
TABLE IV: Ablation study results of different variants of IsoSEL on graph clustering tasks in terms of clustering performance (%).
Cora Citeseer AMAP Photo Pubmed
@@
LSEnet 63.97$\pm$0.67 63.35$\pm$0.56 52.26$\pm$1.09 48.01$\pm$1.25 71.12$\pm$1.30 65.08$\pm$0.73 55.03$\pm$0.79 42.15$\pm$1.02 34.25$\pm$0.24 35.16$\pm$0.30
w/o CIncorp 63.45$\pm$0.92 62.78$\pm$0.88 51.96$\pm$1.45 48.27$\pm$1.65 72.65$\pm$1.92 64.37$\pm$1.35 54.78$\pm$1.03 42.86$\pm$1.62 36.23$\pm$0.50 37.12$\pm$0.69
w/o IsoAug 62.89$\pm$1.05 62.15$\pm$0.97 51.42$\pm$1.58 47.93$\pm$1.80 72.23$\pm$2.01 64.01$\pm$1.50 54.34$\pm$1.12 42.54$\pm$1.79 35.97$\pm$0.64 36.85$\pm$0.82
w/o TRefine 62.34$\pm$1.20 61.78$\pm$1.10 50.87$\pm$1.70 47.42$\pm$1.95 71.89$\pm$2.14 63.78$\pm$1.68 53.98$\pm$1.25 42.21$\pm$1.98 35.78$\pm$0.81 36.64$\pm$0.95
w/o LGroup 61.78$\pm$1.32 61.23$\pm$1.18 50.31$\pm$1.85 46.89$\pm$2.10 71.54$\pm$2.28 63.45$\pm$1.80 53.56$\pm$1.38 41.89$\pm$2.12 35.56$\pm$0.95 36.42$\pm$1.12
IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 74.54$\pm$0.44 69.24$\pm$0.52 55.93$\pm$0.53 43.67$\pm$0.35 36.91$\pm$0.33 38.52$\pm$0.23
TABLE V: Comparison between LSEnet and CSE on Cora, Citerseer, Football and Karate datasets in terms of NMI and ARI (%).
Cora Citeseer Football Karate
NMI ARI NMI ARI NMI ARI NMI ARI
CSE 60.82$\pm$0.30 59.36$\pm$0.17 47.12$\pm$1.01 45.67$\pm$0.67 79.23$\pm$0.12 54.06$\pm$0.05 81.92$\pm$0.29 69.77$\pm$0.02
IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 82.35$\pm$0.53 57.36$\pm$0.35 83.28$\pm$0.72 74.15$\pm$0.55
Performance Gain 3.74 4.66 6.06 4.78 3.12 3.30 1.36 4.38
@@
Figure 2: Parameter sensitivity on dimension of structural entropy.
VIII-C Comparison with Classic Structural Entropy [16]
In this section, we analyze the strength of the proposed IsoSEL against the classic structural entropy [16] (referred to as CSE for short). (1) Time Complexity. We report the running time under different heights of partitioning trees in Fig 2. IsoSEL shows better scalability than CSE. Concretely, CSE has competitive time cost to IsoSEL on the small dataset (Football), but performs undesirably on larger datasets, e.g., at least 13×13$\times$13 × time cost to IsoSEL on Cora. In addition, the linear complexity of IsoSEL is regardless of the height of partitioning tree, while CSE scales exponentially over the height, e.g., $\mathcal{O}(N^{3}\log^{2}N)$ for height $3$, making it prohibited on large graphs. (2) Clustering Effectiveness. Note that, CSE itself neither generates cluster assignment nor node representations. We invoke the heuristic algorithm to construct a partitioning tree of height $2$, where the leaf-parent relationship is regarded as cluster assignment, and thus we instantiate IsoSEL to generate a deep tree of the same height for fair comparison. We summarize the results on Cora, Citeseer, Football and Karate in Table V. IsoSEL consistently outperforms CSE for all the cases. It suggests that cluster assignment is improved by the hyperbolic geometry and attribute incorporation of IsoSEL.
@@ VIII-D Comparison with Our Previous LSEent [20]
This subsection demonstrates the significant improvement compared to LSEnet in the conference version. LSEnet constructs the deep partitioning tree in hyperbolic space, while IsoSEL leverages Lorentz tree contrastive learning to refine the deep tree of LSEnet. First, we compare the clustering effectiveness and report the results on Cora, Citeseer, AMAP, Photo and Pubmed datasets in Table IV. IsoSEL consistently achieves better performance in both NMI and ARI. The performance gain is attributed to hyperbolic tree representation refinement and attribute incorporation via Lorentz tree contrastive learning. Second, we examine the running time under different tree heights in Fig. 2. IsoSEL outperforms LSEnet on both datasets owing to the reduced complexity in the proposed reformulation and the sampling strategy of contrastive learning. In other words, IsoSEL in this paper exhibits superior clustering ability to the previous LSEnet, with minimal computational overhead. (The superiority in expressiveness and imbalanced clusters is elaborated in Sec. VIII-G and Sec. VIII-H, respectively.)
TABLE VI: Ablation study results of different representation space of partitioning tree of IsoSEL.
@@
Hyperbolic 64.56 64.02 53.18 50.45 36.91 38.52
Hyperspherical 50.53 51.23 43.53 40.09 27.35 27.26
Euclidean 56.33 53.78 48.92 46.42 30.07 30.19
@@ VIII-E Ablation Study
This subsection introduce several variants to examine the effectiveness of each proposed module of IsoSEL, and comparison results are summarized in Table IV.
VIII-E1 Contrastive Incorporation vs Input Incorporation
We discuss the strategy for incorporating attribute information into partitioning tree. In IsoSEL, we propose a contrastive approach that incorporates attribute information through mutual information maximization between the original and augmented graphs. Correspondingly, a variant named $\operatorname{w/oCIncorp}$ is introduced to perform input incorporation by directly constructing the partitioning tree from the augmented graph with node attributes. As shown in Table IV, IsoSEL consistently outperforms $\operatorname{w/oCIncorp}$ in all cases. The superiority of contrastive incorporation lies in its ability to not only inject the attribute information but also to refine the tree representations in hyperbolic space, thereby improving cluster assignment.
VIII-E2 Isometric Augmentation vs Weighted-Sum Augmentation
We study how to effectively integrate node attributes into graph structure in the augmented view. In IsoSEL, we derive the isometric augmentation in Eq. (24) which conducts the integration from the representation perspective with the notion of isometry. To evaluate the effectiveness of the proposed formulation, we introduce the $\operatorname{w/oIsoAug}$ variant that generates the augmented structure by a weighted sum of original structure and attribute similarity. In Table IV, we find that the weighted-sum augmentation of $\operatorname{w/oIsoAug}$ leads to inferior performance, demonstrating the effectiveness of isometric augmentation.
VIII-E3 Representation Refinement by Tree Contrastive Learning
This part verifies that contrastive learning itself contributes to the success of clustering. To isolate the impact of attribute information, we introduce the $\operatorname{w/oTRefine}$variant, which is a degenerated version of IsoSEL. Specifically, we replace isometric augmentation with random structure perturbation, where the augmented view is generated by randomly adding or removing edges from the original structure. In Table IV, we find that $\operatorname{w/oTRefine}$ generally achieves a performance gain compared to LSEnet. The contrastive learning refines the partitioning tree representation in hyperbolic space, which is important to clustering performance.
VIII-E4 Neural Lorentz Group for Tree Contrastive Learning
@@
To investigate its effectiveness, we design a variant named $\operatorname{w/oLGroup}$, which replaces proposed Lorentz projection head of neural Lorentz group by the Lorentz layer in [21]. Specifically, the Lorentz layer preserves the manifold structure in the Lorentz model but lacks the linearity of the proposed transformation. We emphasize that linearity ensures consistency between leaf and parent nodes in the partitioning tree, as elaborated in Theorem VII.2. In Table IV, IsoSEL achieves better clustering results than $\operatorname{w/oLGroup}$, demonstrating the importance of consistency in tree contrastive learning.
VIII-E5 Representation Space of Partitioning Tree
@@
We conduct a geometric ablation to study the alignment between hyperbolic space and the partitioning tree. Note that, we are interested in the promising space of the deep tree, while the underlying manifold of input graph is orthogonal to our focus. Theoretically, Theorem VI.1 shows the connection between the tree and hyperbolic space. Empirically, we instantiate the proposed IsoSEL in hyperbolic, hyperspherical and Euclidean space without loss of generality. The clustering results on Cora, Citeseer and Pubmed are given in Table VI. These results show that hyperbolic space is well suited to model the partitioning tree and achieves the best performance among constant curvature spaces.
Figure 3: Sensitivity Analysis of $\alpha$
(a) NMI
(b) ARI
Figure 4: Parameter sensitivity on dimension of structural entropy.
VIII-F Hyperparameter Sensitivity
In this subsection, we discuss the hyperparameter sensitivity of the proposed IsoSEL: attribute augmentation ratio $\alpha$ and height of partitioning tree $H$.
VIII-F1 Attribute Augmentation Ratio
In the proposed isometric augmentation, the hyperparameter $\alpha$ controls the relative importance over attribute and structural information. To evaluate the sensitivity, we vary $\alpha$ in the range of $[0.2,0.8]$, while fixing other hyperparameters. We report the corresponding performance on Cora, Citeseer, AMAP and Computer datasets in Fig. 3. Note that, only slight changes of clustering results is observed under different ratios, and IsoSEL exhibits robustness regarding $\alpha$.
VIII-F2 Height of Partitioning Tree
The hyperparameter $H$ determines the height of the partitioning tree. We collect the clustering results with different tree heights in Fig 4, where $H$ takes the values in $\{2,3,4,5\}$. There are two findings from the empirical results: First, the height of partitioning tree affects the performance of IsoSEL. Second, we notice that IsoSEL achieves the best results on Cora at the height of $2$ while the optimal height of AMAP is $3$. We argue that the optimal height is depend on the inherent complexity of the graph.
VIII-G Embedding Expressiveness of IsoSEL
IsoSEL generates cluster assignment as well as node embedding in hyperbolic space. Here, we evaluate the embedding expressiveness with link prediction. To this end, IsoSEL is compared with the Euclidean models (including GCN [3], SAGE [50], GAT [4] and DGI [51]), hyperbolic models (including HGCN [52] and LGCN [53]), a recent $Q$GCN in ultra hyperbolic space [54] and our previous LSEnet. Popular metrics of AUC (Area Under the Curve) and AP (Average Precision) are employed to quantify the results. The link prediction performances on the five datasets are collected in Table VII, where the height of partitioning tree is set as $3$ for both LSEnet and IsoSEL. As shown in Table IV, Riemannian models generally outperforms the Euclidean counterparts, and IsoSEL achieves the state-of-the-art results. It suggests that Riemannian manifolds better model the complexity of graph structures and, more importantly, the partitioning tree encodes fruitful information for link prediction. Also, compared to LSEnet, IsoSEL consistently obtains the performance gain thanks to the proposed Lorentz tree contrastive learning.
@@ VIII-G Embedding Expressiveness of IsoSEL
GCN 91.19$\pm$0.51 91.85$\pm$0.43 90.16$\pm$0.49 91.91$\pm$0.24 90.12$\pm$0.72 91.08$\pm$0.08 93.86$\pm$0.36 93.24$\pm$0.32 91.16$\pm$0.14 89.96$\pm$0.21
SAGE 86.02$\pm$0.55 88.24$\pm$0.87 88.18$\pm$0.22 89.15$\pm$0.36 98.02$\pm$0.32 97.51$\pm$0.21 92.82$\pm$0.20 93.12$\pm$0.14 89.22$\pm$0.87 89.44$\pm$0.82
GAT 92.55$\pm$0.49 90.27$\pm$0.23 89.32$\pm$0.36 90.08$\pm$0.06 98.67$\pm$0.08 97.21$\pm$0.34 95.93$\pm$0.19 94.57$\pm$0.24 91.60$\pm$0.37 90.37$\pm$0.21
DGI 90.02$\pm$0.80 90.61$\pm$1.00 94.53$\pm$0.40 95.72$\pm$0.10 98.24$\pm$0.13 96.84$\pm$0.24 95.84$\pm$0.05 95.22$\pm$0.36 91.24$\pm$0.60 92.23$\pm$0.50
HGCN 93.60$\pm$0.37 93.64$\pm$0.12 94.39$\pm$0.42 94.38$\pm$0.11 98.06$\pm$0.29 96.84$\pm$0.15 96.88$\pm$0.53 96.18$\pm$0.13 96.30$\pm$0.28 94.72$\pm$0.52
LGCN 92.69$\pm$0.26 91.09$\pm$0.37 93.49$\pm$1.11 94.96$\pm$0.18 97.08$\pm$0.08 96.73$\pm$0.52 96.37$\pm$0.70 95.66$\pm$0.24 95.40$\pm$0.23 93.12$\pm$0.24
$Q$GCN 95.22$\pm$0.29 94.12$\pm$0.51 94.31$\pm$0.73 94.08$\pm$0.34 95.17$\pm$0.45 93.94$\pm$0.18 95.10$\pm$0.03 94.28$\pm$0.33 96.20$\pm$0.34 94.24$\pm$0.31
LSEnet 95.51$\pm$0.60 95.40$\pm$0.04 94.42$\pm$1.51 94.69$\pm$0.05 98.75$\pm$0.67 97.04$\pm$0.13 97.06$\pm$1.02 97.06$\pm$0.05 96.41$\pm$0.10 95.46$\pm$0.10
IsoSEL 96.23$\pm$0.14 95.94$\pm$0.21 95.02$\pm$0.34 95.81$\pm$0.29 98.78$\pm$0.19 97.26$\pm$0.17 97.15$\pm$0.34 97.34$\pm$0.38 96.50$\pm$0.07 95.72$\pm$0.09

### v2

@@ VIII-A1 Datasets & Baselines
VGAE ARGA MVGRL Sublime GRACE DCRN gCooL S3GC DinkNet GC-Flow DGCluster ASIL Avg. Gain
Cora 64.72 68.40 80.03 71.30 73.87 66.91 77.03 74.27 78.11 78.92 76.45 82.14 8.50
Citeseer 61.00 60.72 61.23 68.53 63.12 70.86 71.35 68.80 70.36 70.06 70.14 72.73 5.81
AMAP 74.26 69.28 41.07 47.22 67.66 77.94 68.52 75.15 79.23 72.54 78.92 79.31 10.97
Computer 42.44 45.67 29.53 45.52 46.62 30.42 44.50 39.03 43.52 43.20 43.65 49.03 7.75
Pubmed 68.31 64.82 67.01 66.62 63.72 69.63 68.24 71.30 69.42 73.42 66.39 74.47 6.39
@@ VIII-A3 Model Configuration & Reproducibility
We instantiate the proposed ASIL in the Lorentz model of hyperbolic space with a standard curvature of $\kappa=-1$. The model configuration by default is specified as follows. We consider the partitioning tree of height $2$ by default and hyperparameter $N_{1}$ of network architecture is set as $10$. The augmentation ratio $\gamma$ is $0.01$, the $k$ nearest neighbors for sparsifying $\tilde{\mathbf{A}}$ is set to $8$. We stack the Lorentz convolution once in LSEnet, The parameters are optimized by Adam, the further details and the hyperparameter of the baselines are provided in implementation note of Appendix VI.
@@ VIII-B1 Effectiveness
We report the clustering results in terms of NMI and ARI on the five real datasets in Table I, and ACC in Table II. Deep clustering models require cluster number $K$ in training phase, except DGCluster, RGC and our ASIL. For clustering models without $K$, DGCluster conducts self-supervised representation learning with modularity [3], and RGC searches the optimal $K$ with reinforcement learning [34]. For Riemannian models, i.e., Congregate and ASIL, we project Euclidean attributes onto the respective manifold, and the projection is given by exponential map with the reference point of the origin. DCRN does not scale well to the larger Computer dataset, and runs into Out-Of-Memory (OOM) error on our hardware. In Table I, the proposed ASIL has the best clustering results except two cases, NMI on AMAP and ARI on Computer. Note that, ASIL outperforms $20$ recent deep clustering models by an average of $12.42\%$ NMI on Citeseer, and generally achieves the best ACC in Table II. Also, we observe that ASIL shows more consistent performance over different datasets, e.g., DGCluster obtain the best NMI on AMAP but is unsatisfactory on Computer.
@@ VIII-B2 Efficiency
Table IV collects the theoretical time complexity and empirical running time on Citeseer dataset. The running time is given in seconds, and we are presented as the fastest deep model in the table. As for theoretical complexity, $N$ and $|E|$ denote the number of nodes and edges, respectively. $d$ is the hidden dimension of models. $B$ is the batch size of DinkNet. $S$ is the average degree of nodes and thereby the complexity of S3GC is related to the number of edges. In DGCluster, $P$ is a dimensionality of auxiliary matrix, $P\ll N$. In ASIL, $k$ is the number of nearest neighbors for sparsifying $\tilde{A}$, and it yields the linear complexity thanks to the reformulation of differential structural information and adjacency sparsifying. It is noteworthy to mention that, compared to recent deep clustering model such as S3GC, we achieve superior clustering effectiveness while removing the prior knowledge of cluster number without computational overhead. The following parts will demonstrate how the differential formulation, hyperbolic space and neural Loretnz boost contribute to the success of the proposed model.
TABLE III: Comparison to LSEnet and CSE on Cora, Citerseer, Computer, AMAP and PubMed datasets in terms of NMI and ARI (%).
Cora Citeseer Computer AMAP PubMed
@@
[…]
Average Gain 3.39 3.37 5.32 4.93 4.69 4.44 1.28 3.62 3.88 4.58
TABLE IV: Comparison of time complexity and time cost for baselines.
Method Time Complexity Running Time (s)
DGI $\mathcal{O}(|E|d+Nd^{2})$ 100.7
MVGRL $\mathcal{O}(N^{2}d+Nd^{2})$ 829.8
S3GC $\mathcal{O}(NSd^{2})$ 1247.2
Dink-Net $\mathcal{O}(BKd+K^{2}d+Bd)$ 217.7
DGCluster $\mathcal{O}(K^{2}Nd+P^{2}Nd^{2})$ 79.3
ASIL (Ours) $\mathcal{O}(|E|d+kNd^{2})$ 22.1
VIII-C Comparison with Classic Structural Entropy [27]
In this section, we analyze the strength of the proposed ASIL against the classic structural entropy [27] (referred to as CSE for short), and LSEnet [55]. (1) Time Complexity. We report the running time under different heights of partitioning trees in Fig 2. ASIL shows better scalability than CSE. Concretely, CSE has competitive time cost to ASIL on the small dataset (Football, right image of Fig. 2), but performs undesirably on larger datasets, e.g., an average of $30\times$ time cost to ASIL on Cora (left image of Fig. 2). In addition, the linear complexity of ASIL is regardless of the height of partitioning tree, while CSE scales exponentially over the height, e.g., $\mathcal{O}(N^{3}\log^{2}N)$ for height $3$, making it prohibited on large graphs. (2) Clustering Effectiveness. Note that, CSE itself neither generates cluster assignment nor node representations. We invoke the heuristic algorithm to construct a partitioning tree of height $2$, where the leaf-parent relationship is regarded as cluster assignment, and thus we instantiate ASIL to generate a deep tree of the same height for fair comparison. We summarize the results on Cora, Citeseer, AMAP, Computers and PubMed in Table III. ASIL consistently outperforms CSE and LSEnet for all the cases. It suggests that cluster assignment is improved by the hyperbolic geometry, attribute incorporation and neural Lorentz boost construction of ASIL.
@@ VIII-C Comparison with Classic Structural Entropy [27]
Figure 2: Running time comparison between ASIL, LSEnet and CSE.
@@ VIII-D Comparison with Our Previous LSEent [54]
This subsection demonstrates the significant improvement compared to LSEnet in the conference version. LSEnet constructs the deep partitioning tree in hyperbolic space, while ASIL leverages neural Lorentz boost construction to refine the deep tree of LSEnet. First, we compare the clustering effectiveness and report the results on Cora, Citeseer, AMAP, Photo and Pubmed datasets in Table V. ASIL consistently achieves better performance in both NMI and ARI. The performance gain is attributed to hyperbolic tree representation refinement and attribute incorporation via Neural Lorentz boost construction. Second, we examine the running time under different tree heights in Fig. 2. ASIL outperforms LSEnet on both datasets owing to the reduced complexity in the proposed reformulation and the sparsifying process in the neural Lorentz boost construction. That is, ASIL significantly outperforms LSEnet in both effectiveness and efficiency on the benchmark datasets. (The superiority in expressiveness and imbalanced clusters is elaborated in Sec. VIII-G and Sec. VIII-H, respectively.)
Figure 3: Sensitivity analysis on model hyperparameters $\gamma$, $k$ and $H$.
TABLE V: Ablation study results of different variants of ASIL on graph clustering tasks in terms of clustering performance (%).
Cora Citeseer AMAP Computers PubMed
NMI ARI NMI ARI NMI ARI NMI ARI NMI ARI
[…]
TABLE VI: Ablation study results of different representation space of partitioning tree of ASIL.
@@
Hyperbolic 65.78 64.72 55.01 51.77 38.13 39.74
Hyperspherical 52.39 51.75 44.93 40.77 28.05 28.17
Euclidean 58.39 55.77 49.72 47.92 31.51 31.69
@@ VIII-E Ablation Study
This subsection introduces several variants to examine the effectiveness of each proposed module of ASIL, and comparison results are summarized in Table V.
VIII-E1 Importance of fusing a virtual graph
We examine the performance gain achieved by fusing a virtual graph $\mathbf{A}^{\gamma}$ constructed from hyperbolic node representations. To this end, we introduce a variant that disables the construction of $\tilde{\mathbf{A}}$ and instead uses only the original topology $\mathbf{A}$, which indeed reduces to LSEnet. ASIL significantly outperforms LSEnet on all datasets, particularly on AMAP that with imbalanced classes, where semantic are crucial for distinguishing topological noise. This validates that the inclusion of a virtual enables ASIL to discover clusters enhanced by both structural and semantic perspectives.
VIII-E2 Importance of consistency insurance, achieved by neural Lorentz boost
@@
To investigate the role of consistency insurance in tree contrastive learning, we design a variant named w/oLBoost, which replaces our neural Lorentz group layer with the standard Lorentz linear layer from [9]. While the baseline layer preserves the manifold structure, it lacks the isometric linearity that ensures geometric consistency between leaf and internal nodes in the partitioning tree, as formalized in Theorem VI.3. Table V shows that ASIL achieves significantly better clustering than w/oLBoost, demonstrating that consistency insurance are critical for high-fidelity tree contrastive learning.
VIII-E3 Implicit or explicit tree contrastive learning
In this part, we demonstrate that the implicit tree contrastive learning by a parameterized virtual graph outperforms explicit tree contrastive learning. To this end, we give another variant named as ExplicitTCL. Specifically, we use an fixed isometric augmentation (details in Appendix X) to generate an augmentation view of original graph, and then use neural Lorentz boost to compute tree contrastive loss explicitly as in Eq. 36. Thus, the total loss is given as $\mathcal{H}^{\mathcal{T}_{\text{net}}}(G^{\gamma})+\lambda\mathcal{L}_{\text{TCL}}$, where $\lambda=0.01$ is a coefficient. The results are shown in the third line of Table V. It shows that the proposed implicit approach consistently achieves better performance on the five datasets, which verifies the effectiveness of augmented structural entropy.
VIII-E4 Representation Space of the Partitioning Tree
@@
We conduct a geometric ablation to study the alignment between the underlying manifold and the hierarchical partitioning tree. Note that our focus is on the geometry of the learned tree, not the input graph manifold. Theoretically, Theorem V.1 establishes the low-distortion embeddability of trees in hyperbolic space. Empirically, we instantiate ASIL in three constant-curvature spaces: hyperbolic ($\kappa<0$), hyperspherical ($\kappa>0$), and Euclidean ($\kappa=0$), keeping all other components identical. The clustering results on Cora, Citeseer, and Pubmed are given in Table III. These results confirm that hyperbolic space is uniquely suited for modeling hierarchical partitioning trees, consistently achieving the best performance across datasets.
@@
[…]
VIII-F Hyperparameter Sensitivity
In this subsection, we analyze the sensitivity of ASIL to three key hyperparameters: (i) the number of neighbors $k$ for sparsifying $\tilde{\mathbf{A}}$, (ii) the coefficient $\gamma$ of Lorentz boost construction $\mathbf{A}^{\gamma}=(1-\gamma)\mathbf{A}+\gamma\tilde{\mathbf{A}}$, and (iii) the height $H$ of the partitioning tree (i.e., the dimension of structural entropy).
VIII-F1 Coefficient for Lorentz Boost Construction
The coefficient $\gamma$ controls the importance of representational perspective in the graph fusion. We test $\gamma\in\{0.01,0.05,0.1,0.5,1.0,2.0\}$. When $\gamma=0$, ASIL degenerates to the original LSEnet. As shown in Figure 3(a–e), when $\gamma$ gets larger from $0.5$ to $1.0$, the performance of ASIL gets worse sharply since too much information injection may be treated as noisy structure, especially when $\gamma=2.0$ that exceeds the scope defined in Theorem VII.2.
VIII-F2 $k$-NN Neighborhood Size ($k$)
The virtual graph $\tilde{\mathbf{A}}$ is constructed via $k$-nearest neighbors in Lorentz space. We vary $k\in\{5,8,10,15,30,60\}$ and fix other hyperparameters. As shown in Figure 3(f–j), a small $k$ (e.g., $k=5,8,10$) achieves well performance while a large $k$ (e.g., $k=30,60$) introduces noisy cross-cluster edges, destroying clustering effectiveness.
VIII-F3 Height of Partitioning Tree ($H$)
The hyperparameter $H$ determines the depth of the hierarchical partitioning tree. We evaluate $H\in\{2,3,4,5\}$ and report results in Figure 3(k–o). We observe that performance is sensitive to $H$, with optimal values depending on graph complexity: ASIL achieves peak NMI on Cora at $H=2$, while AMAP prefers $H=3$. This suggests that deeper trees benefit more complex graphs with hierarchical community structures. Notably, ASIL remains stable across a wide range of $H$, demonstrating robustness to model capacity. In summary, ASIL is generally robust to hyperparameter choices, and default values ($k=8$, $\alpha=0.01$, $H=2$) work well across datasets.
VIII-G Embedding Expressiveness of ASIL
ASIL generates cluster assignment as well as node embedding in hyperbolic space. Here, we evaluate the embedding expressiveness with link prediction. To this end, ASIL is compared with the Euclidean models (including GCN [25], SAGE [20], GAT [60] and DGI [61]), hyperbolic models (including HGCN [8] and LGCN [76]), a recent QQGCN in ultra hyperbolic space [67] and our previous LSEnet. Popular metrics of AUC (Area Under the Curve) and AP (Average Precision) are employed to quantify the results. The link prediction performances on the five datasets are collected in Table VII, where the height of partitioning tree is set as $3$ for both LSEnet and ASIL. As shown in Table VII, Riemannian models generally outperforms the Euclidean counterparts, and ASIL achieves the state-of-the-art results. It suggests that Riemannian manifolds better model the complexity of graph structures and, more importantly, the partitioning tree encodes fruitful information for link prediction. Also, compared to LSEnet, ASIL consistently obtains the performance gain thanks to the implicit Lorentz tree contrastive learning.